# Harvard Undergraduate Survey on Generative AI

### v1

@@ 1.1 The Rise of Generative AI: A Student Perspective
In the past few years, generative AI technology — that is, AI systems trained to produce original text, images, and videos — has improved rapidly. In 2019, the state-of-the-art language model was OpenAI’s GPT-2 [10], which could barely form coherent paragraphs. A few years later, ChatGPT surprised many people with its ability to compose poetry, write code, and converse fluently. Today, the most advanced chatbots powered by GPT-4 and Claude 3 can generate stunning images in any art style, perform basic autonomous tasks like browsing the web [8], and consistently outperform most humans on standardized tests. [^1] Claude 3 has an accuracy of 60% on GPQA, a dataset of extremely difficult science questions [1]. For reference, human PhD students score 65% on questions in their field of expertise, and only 34% on questions from a different field — even with access to Google. A recent survey of thousands of AI experts estimated that there is a 50% chance of AI outperforming humans in every possible task by 2047 [7].
@@
If this trend of AI progress continues, generative AI will also have a large impact on the experience of students more broadly. AI may increase mental health problems by introducing uncertainty around career options and challenging the notion of human exceptionalism [6]. Generative AI products aren’t cheap and will likely get more expensive over time (ChatGPT Pro costs \$20 per month), which may exacerbate existing economic inequalities among students.
Finally, AI poses looming questions for society at large. How will we ensure that the companies that develop powerful AI models deploy them responsibly and safely? What will the function of humans be in an economy transformed by automated cognitive labor? What role should an academic institution like Harvard play in influencing crucial AI policy decisions and the development of generative AI more broadly?
In an attempt to shed light on current student perspectives surrounding these questions, the Harvard Undergraduate Association [^2] In collaboration with the AI Safety Student Team, the Harvard Undergraduate Open Data Project, and the Harvard Undergraduate Machine Intelligence Community. commissioned the inaugural “Survey on Generative AI” this April, which was sent to all students enrolled in Harvard College. The survey received 326 responses. In this report, we present key findings from this survey and offer recommendations and perspectives based on the results.

### v2

@@ 1.1 The Rise of Generative AI: A Student Perspective
In recent years, generative AI technology — AI systems trained to produce original text, images, and videos — has advanced rapidly. Today, the most advanced chatbots powered by GPT-4 and Claude 3 can generate stunning images in any art style, perform basic autonomous tasks like browsing the web [8], and consistently outperform most humans on standardized tests. [^1] Claude 3 has an accuracy of 60% on GPQA, a dataset of extremely difficult science questions [1]. For reference, human PhD students score 65% on questions in their field of expertise, and only 34% on questions from a different field — even with access to Google. A recent survey of thousands of AI experts estimated that there is a 50% chance of AI outperforming humans in all tasks by 2047 [7].
@@
If this trend of AI progress continues, generative AI will also have a large impact on the experience of students more broadly. AI may affect mental health [6] and economic inequality among students. Moreover, AI raises important societal questions about responsible deployment and the future role of humans in an AI-transformed economy.
In an attempt to shed light on current student perspectives surrounding these questions, the Harvard Undergraduate Association [^2] In collaboration with the AI Safety Student Team, the Harvard Undergraduate Open Data Project, and the Harvard Undergraduate Machine Intelligence Community. commissioned the inaugural “Survey on Generative AI” this April. In this report, we present key findings: use of generative AI among students is nearly pervasive, income levels shape the impact of AI usage, and students have a wide range of opinions about AI’s future, reflecting both optimism and concern.